OpenStudy (anonymous):

Calulate the slope of the line between points (-3,-9) and (5,-7) -1/4 1/4 4 -4

OpenStudy (anonymous):

(-3,-9) and (5,-7) m = (-7 + 9)/ ( 5 + 3) = 2/8 = 1/4

OpenStudy (anonymous):

use this. .

Directrix (directrix):

1/4 -2/-8 = 1/4

OpenStudy (ash2326):

@lisa69 did you understand?

OpenStudy (katrinakaif):

Your given points: (-3,-9) and (5,-7) (-3,-9) : (x1,y1) (5,7) : (x2,y2) To find the slope of the given points, you use the equation: y2 - y1 ------ x2 - x1 Now we simply substitute: 7 - (-9) ------ 5 - (-3) Then simplify in order to conclude to your answer

OpenStudy (anonymous):

Yes a little I'm new at this

OpenStudy (ash2326):

Check @KatrinaKaif 's reply, she has illustrated it beautifully

OpenStudy (katrinakaif):

@ash2326 Danke :-)

OpenStudy (pokemon23):

@KatrinaKaif YOUR AWESOME

OpenStudy (anonymous):

all such a happy family :)

OpenStudy (katrinakaif):

@pokemon23 Thankyou, you are as well ;-)

OpenStudy (anonymous):

Well, slope means "rise over run"

OpenStudy (anonymous):

@KatrinaKaif: I think y2 - y1 ------ x2 - x1 is not a equation ...

OpenStudy (pokemon23):

your right ffm

OpenStudy (katrinakaif):

Formula; Beg my pardon

OpenStudy (anonymous):

Okay now medal for you! :)